OpenStudy (anonymous):

can some one help me with six home work problems here so lost doing derivatives.

OpenStudy (anonymous):

Thats what I am working on tonight, what's your problems? Also if you haven't checked it out Khan Academy.com has lots of great videos

OpenStudy (anonymous):

the book wants me to differentiate the function g(t)=2t ^ -3/4

OpenStudy (anonymous):

I tried solving this but my answer is way to far off from the book

OpenStudy (anonymous):

Did you use the exponent rule?

OpenStudy (anonymous):

yes i did, sorry my pc froze

OpenStudy (anonymous):

f'=n(c)x^n-1 2(-3/4)t^-3/4-1 -1.5t^-7/4? that what you got?

OpenStudy (anonymous):

yeah it's -3/4 t ^-7/4

OpenStudy (anonymous):

sorry -3/2t ^ 7/4

OpenStudy (anonymous):

Hey cool, your boosting my confidence here, LOL I got a test tomorrow night

OpenStudy (anonymous):

I got time for one more before I get to get to bed, whats another?

OpenStudy (anonymous):

A(s)= -12/s^5

OpenStudy (anonymous):

I think you need to rewrite it first A(s)=-12*s^-5 then use the exponent rule rule again so -60*s^-6 Then move it back down -60/s^6 that what you got?

OpenStudy (anonymous):

yeah you are good at this :)

OpenStudy (anonymous):

It was actually one of those that brought me to this web site tonight

OpenStudy (anonymous):

Are you just working on the exponent rule?

OpenStudy (anonymous):

ah ha

OpenStudy (anonymous):

yeah there are other one's with radicals

OpenStudy (anonymous):

yeah those aren't much fun

OpenStudy (anonymous):

yes headache

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

ha its grtting really late over here S(s)= \[\sqrt{p}- p\]

OpenStudy (anonymous):

Oh all mine so far have been in the demonstrator first get rid of the root it is p^1/2-p

OpenStudy (anonymous):

then, get rid of the constant and use the exp rule so 1/2p^-1/2

OpenStudy (anonymous):

Sorry it is getting late and figure out that equation thing isn't fast enough lol

OpenStudy (anonymous):

yeah the answer was 1/2p-1/2-1 ?

OpenStudy (anonymous):

That what the book says?

OpenStudy (anonymous):

yeah 1/2 p^-1/2 -1

OpenStudy (anonymous):

lol tired, Well good luck with the rest Night

OpenStudy (anonymous):

yeah thank you so much :)

OpenStudy (anonymous):

welcome